OpenStudy (anonymous):

The vertex of this parabola is at (-5, -2). When the x-value is -4, the y-value is 2. What is the coefficient of the squared expression in the parabola's equation?

OpenStudy (cwrw238):

in vertex the parabola is y = a(x -b)^2 + c vertex is at (-5,-2) s0 y = a(x +5)^2 - 2 x = -4 when y = 2 so 2 = a(-4+5^)2 - 2 2 = a - 2 a = 4 therefor 4 is the coefficient of x^2 in the parabola's equation